# Revisiting and Maximizing Temporal Knowledge in Semi-supervised Semantic Segmentation

### v1

@@ Abstract
In semi-supervised semantic segmentation, the Mean Teacher- and co-training-based approaches are employed to mitigate confirmation bias and coupling problems. However, despite their high performance, these approaches frequently involve complex training pipelines and a substantial computational burden, limiting the scalability and compatibility of these methods. In this paper, we propose a PrevMatch framework that effectively mitigates the aforementioned limitations by maximizing the utilization of the temporal knowledge obtained during the training process. The PrevMatch framework relies on two core strategies: (1) we reconsider the use of temporal knowledge and thus directly utilize previous models obtained during training to generate additional pseudo-label guidance, referred to as previous guidance. (2) we design a highly randomized ensemble strategy to maximize the effectiveness of the previous guidance. Experimental results on four benchmark semantic segmentation datasets confirm that the proposed method consistently outperforms existing methods across various evaluation protocols. In particular, with DeepLabV3+ and ResNet-101 network settings, PrevMatch outperforms the existing state-of-the-art method, Diverse Co-training, by +1.6 mIoU on Pascal VOC with only 92 annotated images, while achieving 2.4 times faster training. Furthermore, the results indicate that PrevMatch induces stable optimization, particularly in benefiting classes that exhibit poor performance. Code is available at https://github.com/wooseok-shin/PrevMatch.
Index Terms:
Consistency regularization, semantic segmentation, semi-supervised learning, temporal knowledge
@@ 1 Introduction
Semantic segmentation [1, 2, 3, 4] is a critical task in various computer vision-related applications, such as autonomous driving [5, 6, 7], medical image analysis [8, 9], robotics [10], among others [11, 12, 13], where the goal is to assign a semantic class label to each pixel in an image. Despite the recent success of supervised learning-based methods in semantic segmentation, obtaining precise pixel-level annotations for supervised learning is extremely time-consuming and expensive. This limits the applicability of supervised learning-based methods across various domains and fields. Thus, extensive research has been conducted in semi-supervised semantic segmentation to overcome this limitation. The research has focused on developing methods enabling models to learn effectively from a limited number of labeled images along with a large number of unlabeled images.
@@
In semi-supervised learning, self-training [14, 15, 16, 17, 18, 19, 20, 21] and consistency regularization [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] have become the predominant approaches for utilizing unlabeled data. In particular, self-training involves generating pseudo-labels using the predictions of the current model at each iteration for unlabeled samples and leveraging them to train the model in conjunction with labeled data. Consistency regularization encourages a network to predict consistently for various perturbed forms of identical input. Recent studies have focused on designing frameworks that combine self-training and consistency regularization to exploit the strengths of each method. However, self-training-based methods still suffer from the confirmation bias problem [17] even when combined with consistency regularization.
@@
This problem is attributed to the accumulation of pseudo-label errors produced by the model itself, which exacerbate as self-training progresses. To mitigate this problem, existing methods distinguish the model prediction processes for supervised outputs and pseudo-labels. In other words, the supervised outputs and pseudo-labels are obtained from different predictions, respectively. Among them, some studies [34, 35, 32] have adopted the weak-to-strong consistency paradigm from the perspective of input separation, which supervises the prediction from a strongly perturbed input using the pseudo-label generated from its weakly perturbed counterpart (Fig. 1a). Its success is based on the idea that more reliable pseudo-labels can be derived from weak perturbations, and strong perturbations aid in mitigating confirmation bias [17, 21], broadening the knowledge and unlabeled data space [36, 32, 29], and shifting the model’s decision boundary to low-density regions [37]. Because of this advantage, the weak-to-strong consistency paradigm has become a fundamental component of the most recent methods.
@@
Figure 1: Illustration of the frameworks for (a) FixMatch [34] (or PseudoSeg [35]), (b) Mean Teacher-based structure [26, 38, 39, 40], (c) Co-training [28, 29] (cps: cross pseudo supervision), (d) Dual Mean Teacher [30, 31], and (e) the proposed method. In (d), the inputs ($x^{s_{cut}},x^{s_{class}}$) indicate the CutMix [41] and ClassMix [42] augmentations used in Dual Teacher [31].
@@
Another method to obtain different prediction views involves using network perturbation based on a teacher-student structure. In particular, Mean Teacher [26] is a representative approach in semi-supervised segmentation, where a teacher network is derived using an exponential moving average (EMA) of the student model’s weights (Fig. 1b). Although the Mean Teacher generates somewhat different prediction views between the teacher and student, this method is limited by a coupling problem [27]: as training progresses, the teacher and student become tightly linked, and consequently, the teacher’s predictions become similar to those of the student. To mitigate the coupling problem and obtain diverse predictions, certain studies [30, 31] have proposed a dual EMA teacher-based framework where two teachers are alternatively updated at every epoch (Fig. 1d). Among them, PS-MT [30] produces more reliable pseudo-labels by ensembling the predictions of the two teachers. By contrast, Na et al. [31] reported that ensembling predictions can reduce the diversity of pseudo-labels. Thus, they proposed a Dual Teacher framework that alternately activates two teachers in each epoch to generate diversified pseudo-labels. Although these studies have demonstrated the benefits of using multiple teachers to mitigate the coupling problem, their ability to provide reliable and diverse pseudo-labels simultaneously remains limited. In addition, they have incorporated additional complex components to ensure diversity between the two teachers. In particular, adversarial feature perturbation and a new loss function are used for PS-MT. Further, distinct types of augmentation provided to each teacher (e.g., CutMix or ClassMix) and layer perturbation are used for Dual Teacher. Consequently, this complexity can hinder the scalability and compatibility of these approaches with existing semi-supervised methods.
@@
Instead of using EMA-based teachers, a co-training paradigm [43, 44, 45, 46, 27, 28, 47, 29] has been widely used to expand prediction views (Fig. 1c). This paradigm involves simultaneously training multiple networks with different initializations in a mutual teaching manner, where each network supervises the others using pseudo-labels generated from its predictions. Building on this concept, subsequent studies [45, 29] have demonstrated that increasing the diversity of pseudo-label views from student networks improves their generalization ability. For example, this can involve the use of more co-networks, diverse input domains (e.g., RGB and frequency), or different architectures (e.g., CNN and Transformer). Co-training provides diverse pseudo-label guidance with stability and without concerns regarding the coupling problem. However, its scalability remains constrained due to computational complexity and resource demands.
@@
In this paper, we propose the PrevMatch framework, which efficiently expands pseudo-label views by maximizing the utilization of previous models obtained during training, as depicted in Fig. 1e. The PrevMatch framework is based on two main ideas. First, to efficiently address the coupling problem, we revisit the utilization of temporal knowledge. Specifically, we save several models at specific epochs during training and utilize their predictions as additional guidance, referred to as previous guidance, which acts as a regularizer in conjunction with standard guidance. This strategy addresses the coupling problem and reduces the complexity associated with additional training components. Second, we design a highly randomized ensemble strategy to maximize the effectiveness of utilizing the previous guidance. This approach involves selecting a random number of models from those previously saved and ensembling their predictions using randomized weights. This strategy can efficiently provide diverse and reliable pseudo-labels, while avoiding the significant computational complexity inherent in co-training approaches.
Extensive experiments conducted across various evaluation protocols on the PASCAL, Cityscapes, COCO, and ADE20K datasets reveal that the proposed PrevMatch significantly outperforms existing methods. In particular, compared to Diverse Co-training, the current state-of-the-art method, PrevMatch achieves a +1.6 mIoU improvement on Pascal VOC with 92 labels while accelerating training by 2.4 times. In addition to quantitative evaluations, ablation studies and analyses of the proposed components are conducted to explain the success of PrevMatch.
@@ 2 Related Work
2.1 Semi-supervised Learning
The central challenge in semi-supervised learning lies in extracting additional supervision for unlabeled samples. Previous studies can be categorized into two main approaches: self-training and consistency regularization. Self-training-based methods [14, 15, 16, 17, 18, 19, 20] assign pseudo-labels to unlabeled data using the ongoing model predictions and integrate them with labeled data to retrain the model. Consistency regularization methods [22, 23, 24, 25, 34, 48, 26, 43, 27, 49] enforce a model to produce invariant predictions for different perturbations, such as input and network perturbations.
In terms of input perturbations [22, 23, 24, 25, 34, 48], the $\Pi$-model [50, 22, 23] applies simple stochastic perturbations, while VAT [24] applies adversarial perturbations to unlabeled samples. Subsequently, UDA [25] expands the perturbation pool using RandAugment [51] to provide diverse augmented samples. Furthermore, FixMatch [34], a standard approach in recent semi-supervised learning, adopts a weak-to-strong paradigm where the prediction for a strongly perturbed input is supervised by pseudo-labels generated from the prediction for a weakly perturbed input. Network perturbation, another approach to enforcing consistency regularization, has also been widely investigated [23, 26, 43, 27, 49]. Temporal Ensembling [23] and Mean Teacher [26] methods leverage earlier predictions and weights of the model, respectively, to obtain different prediction views. In addition, co-training-based methods [43, 27, 49] adopt multiple networks with different initializations and enforce consistency regularization between their predictions. Among the numerous studies discussed, the weak-to-strong paradigm, Mean Teacher structure, and co-training approaches have been widely used to advance semi-supervised semantic segmentation.
@@ 2.2 Semi-supervised Semantic Segmentation
Owing to the detailed nature of the segmentation task, which involves numerous pixels and multiple classes within an image, semantic segmentation requires more time-consuming and costly efforts to annotate labels and demands sophisticated approaches for accurate performance. Thus, recent research has increasingly focused on developing methods for semi-supervised semantic segmentation, based on previous achievements in semi-supervised learning. Early studies [52, 53] employed a GAN framework [54] to generate additional supervision for unlabeled images by distinguishing between pseudo-labels and manual labels. Recent studies have focused on developing improved methods based on weak-to-strong consistency, Mean Teacher, and co-training paradigms. PseudoSeg [35] adopts a weak-to-strong paradigm based on a single network. CPS [28, 45] and GCT [55] employ two networks for co-training and demonstrate that co-training outperforms the Mean Teacher approach. Subsequent studies have integrated the weak-to-strong paradigm (i.e., input perturbation) into the Mean Teacher or co-training approaches (i.e., network perturbation), demonstrating performance improvements.
A stream of research that combines the weak-to-strong paradigm with the Mean Teacher structure has proposed techniques such as advanced data augmentation [38, 56, 39, 40], prototype learning [57], curriculum learning [58], symbolic reasoning [59], and dual Mean Teacher [30, 31]. Among them, PS-MT [30] and Dual Teacher [31] methods implement a dual EMA teacher-based framework to mitigate the coupling problem between the teacher and student models, with two teachers alternately updating each epoch based on the EMA of the student’s weights. However, their pipelines inevitably involve complex components to resolve the coupling problem, limiting the scalability and compatibility of these approaches. Moreover, regarding pseudo-label generation, PS-MT ensembles the predictions of two teachers to improve reliability, whereas Dual Teacher alternately uses the predictions of each teacher to enhance diversity. However, neither method satisfies both reliability and diversity. In other words, PS-MT improves the reliability of pseudo-labels but lacks diversity, whereas Dual Teacher suffers from the opposite limitation.
Another stream of research, combining the weak-to-strong paradigm with co-training, has developed techniques such as a shared backbone with multiple heads [46], conservative-progressive learning [60], and increasing the diversity of co-training [29]. Among them, Li et al. [29] investigated the working mechanism of co-training and discovered that providing distinct pseudo-label views improves generalization ability. To this end, they proposed Diverse Co-training, which incorporates variations in the input domains (RGB and frequency) and architectures (CNN and Transformer). However, in diverse co-training, the number of whole networks to train increases as pseudo-label views increase.
@@
In this study, instead of using Mean Teacher or co-training approaches, we revisit the utilization of temporal knowledge and efficiently expand pseudo-label views by maximizing the use of previous models. The proposed PrevMatch framework can be seen as simplifying and extending [30, 31] while also enhancing the efficiency of [29]. Specifically, we eliminate the complex components used in [30, 31], such as distinct augmentation types for each epoch, layer/adversarial feature perturbations, EMA teachers, and a new loss function. Instead, PrevMatch reuses the previous models and the weakly perturbed input used in the standard flow, thereby improving the simplicity and compatibility of the overall framework. Furthermore, we provide reliable and diverse pseudo-labels to the student network through a highly randomized ensemble strategy. Moreover, although both PrevMatch and co-training provide diverse pseudo-label views, PrevMatch operates on a single trainable network with fixed previous models. This facilitates greater efficiency in terms of computational and memory costs.
@@ 2.3 Temporal Knowledge in Semi-supervised Learning
In the context of semi-supervised learning, several studies have leveraged temporal knowledge obtained from previous training stages. Temporal Ensembling [23] accumulates predictions for unlabeled samples across different epochs using the EMA approach and enforces consistency between the current and EMA predictions. Mean Teacher [26] averages the model weights across the training steps using the EMA approach, producing a more stable teacher network. As a further extension, a dual EMA teacher-based framework was proposed [30, 31], where two teachers are alternately updated at each epoch. Moreover, TC-SSL [20] measures the time-consistency (TC) scores of individual samples across training epochs and selects unlabeled samples with higher TC scores for consistency learning. This approach assumes that time-consistent predictions are typically accurate. Similar to TC-SSL, ST++ [21] proposes a selective re-training scheme that selects more reliable samples based on a stability score. The stability score based on the mIoU metric is derived from comparisons between the predictions of the final model and those of several checkpoints obtained during training. In addition to semi-supervised learning, Feng et al. [61] proposed a temporal consistency framework that adopts two temporary teachers to learn instance temporal consistency in representation learning.
@@ 3.1 Preliminaries & Overall Workflow
Figure 2: Overall framework of PrevMatch (based on UniMatch). FP indicates the feature perturbation used in UniMatch. $g$ and $h$ are the encoder and decoder of the entire network $f$, respectively. When the maximum length of the previous list is exceeded, the oldest model is replaced with the new model using a first-in-first-out (FIFO) approach. The red and blue arrows denote standard and proposed guidance flows, respectively.
  1. 1.
@@
    $L_{s}=\sum_{i,j}H(y^{l}_{ij},p^{l}_{ij}),$ (2)
    where $i$ and $j$ indicate pixel indices.
  2. 2.
    Unsupervised – Standard Flow Following the success of the weak-to-strong consistency paradigm popularized by FixMatch [34], most semi-supervised segmentation methods have adopted this paradigm. Concretely, two perturbed images (i.e., $x^{w}$ and $x^{s}$) are obtained by applying weak and strong augmentations to an unlabeled image $x^{u}$. The network $f$ receives the two perturbed images and outputs the predicted class distributions $p^{w}$ and $p^{s}$. As there is no ground-truth label for the unlabeled image, a pseudo-label for the weakly perturbed prediction is obtained by: $y^{w}\text{= argmax}(p^{w})$. Then, this pseudo-label is used to supervise the strongly perturbed prediction, as depicted in the standard flow presented in Fig. 1e (red arrow). This consistency term can be formulated as follows.
    $C(p^{w},p^{s})=\sum_{i,j}\mathbb{1}(max(p^{w}_{ij})\geq\tau)H(y^{w}_{ij},p^{s}_{ij}),$ (3)
    where $\mathbb{1}(\cdot)$ is an indicator function, and $\tau$ is a confidence threshold used to ignore noise in a pseudo-label. Furthermore, the UniMatch framework [32] introduced two contributions based on FixMatch: the dual stream image-level perturbation and the feature-level perturbation strategies, to expand the perturbation space, thereby achieving significant improvements. Therefore, we select the UniMatch framework as our baseline. In particular, two strongly perturbed images ($x^{s_{1}}$ and $x^{s_{2}}$) are derived by randomly applying a strong augmentation pool to the same image, $x^{u}$. Subsequently, their corresponding predictions ($p^{s_{1}}$ and $p^{s_{2}}$) are generated through the network $f$. Regarding the feature-level perturbation strategy, the prediction is obtained as follows: $p^{fp}=h(\text{Dropout}(g(x^{w})))$, where $g$ and $h$ are the encoder and decoder of the entire network $f$, respectively. Ultimately, three predictions are simultaneously supervised by a common pseudo-label derived from a weak view (red arrow in Fig. 2). This can be formulated as follows.
    $L_{u(standard)}=C(p^{w},p^{s_{1}})+C(p^{w},p^{s_{2}})+C(p^{w},p^{fp}).$ (4)
  3. 3.
    Unsupervised – Proposed Flow To obtain additional pseudo-label guidance, the same weakly perturbed image, $x^{w}$, used in the standard branch, is fed into the previous model branch. Subsequently, $k$ models, where $k$ is chosen randomly from $\{1,2,...,K\}$, are selected from the saved list of previous models, and $k$ predictions for $x^{w}$ are generated. Previous guidance, $p^{w^{\prime}}$, is obtained by aggregating the $k$ predictions using randomized ensemble weights to improve the diversity of the pseudo-label view. As illustrated by the blue arrow in Fig. 2, the previous guidance functions as additional regularizers, supervising the three predictions. This can be formulated as follows.
    $L_{u(prev)}=C(p^{w^{\prime}},p^{s_{1}})+C(p^{w^{\prime}},p^{s_{2}})+C(p^{w^{\prime}},p^{fp}).$ (5)
@@
3.2 Previous Guidance: Revisiting Temporal Knowledge
To address the coupling problem between student and teacher networks, we revisit the utilization of temporal knowledge that can be obtained during the training process. In the literature, Temporal Ensembling [23] and Mean Teacher [26] methods average the temporal knowledge of a model in terms of its predictions and weights, respectively. PS-MT [30] and Dual Teacher [31], inheriting the spirit of the Mean Teacher [26] method, aim to further exploit temporal knowledge by adopting two EMA teachers. TC-SSL [20] and ST++ [21] methods implement a filtering mechanism to exclude less-informative unlabeled samples using specific scoring criteria measuring temporal consistency.

### v2

# PrevMatch: Revisiting and Maximizing Temporal Knowledge in Semi-Supervised Semantic Segmentation

@@ Abstract
In semi-supervised semantic segmentation, the Mean Teacher- and co-training-based approaches are employed to mitigate confirmation bias and coupling problems. However, despite their high performance, these approaches frequently involve complex training pipelines and a substantial computational burden, limiting the scalability and compatibility of these methods. In this paper, we propose a PrevMatch framework that effectively mitigates the aforementioned limitations by maximizing the utilization of the temporal knowledge obtained during the training process. The PrevMatch framework relies on two core strategies: (1) we reconsider the use of temporal knowledge and thus directly utilize previous models obtained during training to generate additional pseudo-label guidance, referred to as previous guidance. (2) we design a highly randomized ensemble strategy to maximize the effectiveness of the previous guidance. PrevMatch, a simple yet effective plug-in method, can be seamlessly integrated into existing semi-supervised learning frameworks with minimal computational overhead. Experimental results on three benchmark semantic segmentation datasets show that incorporating PrevMatch into existing methods significantly improves their performance. Furthermore, our analysis indicates that PrevMatch facilitates stable optimization during training, resulting in improved generalization performance.
@@ 1 Introduction
Semantic segmentation is a critical task in various computer vision-related applications, such as autonomous driving [wang2020deep, xie2021segformer], medical image analysis [fan2020pranet, shin2022comma], and robotics [milioto2019bonnet], where the goal is to assign a semantic class label to each pixel in an image. Despite the recent success of supervised learning-based methods in semantic segmentation, obtaining precise pixel-level annotations for supervised learning is extremely time-consuming and expensive. This limits the applicability of supervised learning-based methods across various domains and fields. Thus, extensive research has been conducted in semi-supervised semantic segmentation to overcome this limitation. The research has focused on developing methods enabling models to learn effectively from a limited number of labeled images along with a large number of unlabeled images.
@@
In semi-supervised learning, self-training [lee2013pseudo, arazo2020pseudo, chen2022debiased, zhou2020time, yang2022st++] and consistency regularization [sajjadi2016regularization, laine2016temporal, miyato2018virtual, xie2020unsupervised, chen2021semi, yang2023revisiting] have become the predominant approaches for utilizing unlabeled data. In particular, self-training involves generating pseudo-labels using the predictions of the current model at each iteration for unlabeled samples and leveraging them to train the model in conjunction with labeled data. Consistency regularization encourages a network to predict consistently for various perturbed forms of identical input. Recent studies have focused on designing frameworks that combine self-training and consistency regularization to exploit the strengths of each method. However, self-training-based methods still suffer from the confirmation bias problem [arazo2020pseudo], where the model becomes overfitted in the wrong direction, even when combined with consistency regularization.
@@
This problem is attributed to the accumulation of pseudo-label errors produced by the model itself, which are exacerbated as self-training progresses. To mitigate this problem, existing methods distinguish the model prediction processes for supervised outputs and pseudo-labels. In other words, the supervised outputs and pseudo-labels are obtained from different predictions, respectively. Some studies [sohn2020fixmatch, zou2020pseudoseg, zhong2021pixel, yang2023revisiting] have adopted the weak-to-strong consistency paradigm from the perspective of input separation, which supervises the prediction from a strongly perturbed input using the pseudo-label generated from its weakly perturbed counterpart (Fig. 1a).
@@
Figure 1: Illustration of the frameworks for (a) FixMatch [sohn2020fixmatch], (b) Mean Teacher-based structure [tarvainen2017mean], (c) Co-training [chen2021semi, li2023diverse] (cps: cross pseudo supervision), (d) Dual Mean Teacher [liu2022perturbed, na2023switching], and (e) the proposed method. In (d), the inputs ($x^{s_{cut}},x^{s_{class}}$) indicate the CutMix [yun2019cutmix] and ClassMix [olsson2021classmix] augmentations used in Dual Teacher [na2023switching].
@@
Another method to obtain different prediction views involves using network perturbation based on a teacher-student structure [laine2016temporal, tarvainen2017mean, jin2022semi, wang2022semi, wang2023hunting]. In particular, Mean Teacher [tarvainen2017mean] is a representative approach in semi-supervised segmentation, where a teacher network is derived using an exponential moving average (EMA) of the student model’s weights (Fig. 1b). Although the Mean Teacher generates somewhat different prediction views between the teacher and student, this method is limited by a coupling problem [ke2019dual]: as training progresses, the teacher and student become tightly linked, and consequently, the teacher’s predictions become similar to those of the student. To mitigate the coupling problem and obtain diverse prediction views, certain studies [liu2022perturbed, na2023switching] have proposed a dual EMA teacher-based framework where two teachers are alternatively updated at every epoch (Fig. 1d). Among them, PS-MT [liu2022perturbed] produces more reliable pseudo-labels by ensembling the predictions of the two teachers. By contrast, Na et al. [na2023switching] reported that ensembling predictions can reduce the diversity of pseudo-labels. Thus, they proposed the Dual Teacher framework that alternately activates two teachers in each epoch to provide diversified pseudo-labels. Although these studies have demonstrated the benefits of using multiple teachers to mitigate the coupling problem, their ability to provide reliable and diverse pseudo-labels over the course of training remains limited, as they typically focus on only one aspect. In other words, PS-MT improves the reliability of pseudo-labels but lacks diversity, whereas Dual Teacher suffers from the opposite limitation. In addition, they have incorporated additional complex components to ensure diversity between the two teachers. In particular, adversarial feature perturbation and a new loss function are used for PS-MT. Further, distinct types of augmentation provided to each teacher (e.g., CutMix or ClassMix) and layer perturbation are used for Dual Teacher. Consequently, this complexity can hinder the scalability and compatibility of these approaches with existing semi-supervised segmentation methods.
@@
Instead of using EMA-based teachers, a co-training paradigm [qiao2018deep, ouali2020semi, fan2022ucc, ke2019dual, chen2021semi, wang2023conflict] has been widely used to expand prediction views (Fig. 1c). This paradigm involves simultaneously training multiple networks with different initializations in a mutual teaching manner, where each network supervises the others using pseudo-labels generated from its predictions. Building on this concept, a subsequent study [li2023diverse] has demonstrated that increasing the diversity of pseudo-label views from student networks improves their generalization ability. Co-training provides diverse pseudo-label guidance with stability and without concerns regarding the coupling problem. However, its scalability remains constrained due to computational complexity and resource demands.
@@
In this paper, we propose the PrevMatch framework, which efficiently expands pseudo-label views by maximizing the utilization of previous models obtained during training, as depicted in Fig. 1e. The PrevMatch framework is based on two main ideas. First, to efficiently address the coupling problem, we revisit the utilization of temporal knowledge. Specifically, we save several models at specific epochs during training and utilize their predictions as additional guidance, referred to as previous guidance, which acts as a regularizer in conjunction with standard guidance. Second, we design a highly randomized ensemble strategy to maximize the effectiveness of utilizing the previous guidance. This approach involves selecting a random number of models from those previously saved and ensembling their predictions using randomized weights. These strategies can efficiently provide diverse and reliable pseudo-labels while avoiding the complexities inherent in dual EMA- and co-training-based approaches. Extensive experiments conducted across various evaluation protocols on the Pascal, Cityscapes, and ADE20K datasets reveal that the proposed PrevMatch significantly improves the performance of different baseline methods. In particular, PrevMatch achieves a 0.9-5.4% mean IoU improvement on Pascal VOC with a 92-label setting.
@@ 2.1 Semi-supervised Semantic Segmentation
Recent studies have focused on developing improved methods based on weak-to-strong consistency, Mean Teacher, and co-training paradigms. PseudoSeg [zou2020pseudoseg] adopts a weak-to-strong paradigm based on a single network. CPS [chen2021semi] and GCT [ke2020guided] employ two networks for co-training and demonstrate that co-training outperforms the Mean Teacher approach. UniMatch [yang2023revisiting] introduces two key enhancements to FixMatch: dual-stream image-level perturbation and feature-level perturbation strategies. Subsequent studies have integrated the weak-to-strong paradigm (i.e., input perturbation) into the Mean Teacher or co-training approaches (i.e., network perturbation), demonstrating performance improvements. A stream of research that combines the weak-to-strong paradigm with the Mean Teacher structure has proposed techniques such as advanced data augmentation [hu2021semi, zhao2023augmentation, zhao2023instance], prototype learning [xu2022semi], curriculum learning [ma2023enhanced], symbolic reasoning [liang2023logic], and dual Mean Teacher [liu2022perturbed, na2023switching]. Another stream of research, combining the weak-to-strong paradigm with co-training, has developed techniques such as a shared backbone with multiple heads [fan2022ucc], conservative-progressive learning [fan2023conservative], and increasing the diversity of co-training [li2023diverse].
@@
The proposed PrevMatch framework can be seen as simplifying and extending PS-MT [liu2022perturbed] and Dual Teacher [na2023switching] while also enhancing the efficiency of Diverse Co-training [li2023diverse]. Specifically, we eliminate the complex components used in PS-MT and Dual Teacher, such as distinct augmentation types for each epoch, layer/adversarial feature perturbations, EMA teachers, and a new loss function. Instead, PrevMatch reuses the previous models and the weakly perturbed input used in the standard flow, thereby improving the simplicity and compatibility of the overall framework. Furthermore, we provide reliable and diverse pseudo-labels to the student network through a highly randomized ensemble strategy. Moreover, although both PrevMatch and Diverse Co-training provide diverse pseudo-label views, PrevMatch operates on a single trainable network with fixed previous models. This facilitates greater efficiency in terms of computational and memory costs.
@@ 2.2 Temporal Knowledge in Semi-supervised Learning
In the context of semi-supervised learning, several studies have leveraged temporal knowledge obtained from previous training stages. Temporal Ensembling [laine2016temporal] accumulates predictions for unlabeled samples across different epochs using the EMA approach and enforces consistency between the current and EMA predictions. Mean Teacher [tarvainen2017mean] averages the model weights across the training steps using the EMA approach, producing a more stable teacher network. PS-MT [liu2022perturbed] and Dual Teacher [na2023switching], inheriting the spirit of the Mean Teacher method, aim to further exploit temporal knowledge by adopting two EMA teachers. Moreover, TC-SSL [zhou2020time] and ST++ [yang2022st++] methods implement a filtering mechanism to exclude less-informative unlabeled samples using specific scoring criteria measuring temporal consistency. In contrast to existing methods, we directly utilize previous models and generate pseudo-label guidance from their predictions without relying on a complex pipeline of dual EMA-based methods or a filtering mechanism.
@@ 3.1 Preliminaries & Overall Workflow
  1. 1)
@@
    $L_{s}=\sum H(y^{l},p^{l}).$ (2)
  2. 2)
    Unsupervised – Standard Flow Following the success of the weak-to-strong consistency paradigm popularized by FixMatch [sohn2020fixmatch], most semi-supervised segmentation methods have adopted this paradigm. Concretely, two perturbed images ($x^{w}$ and $x^{s}$) are obtained by applying weak and strong augmentations to an unlabeled image $x^{u}$. The network $f$ receives the two perturbed images and outputs the predicted class distributions $p^{w}$ and $p^{s}$. As there is no ground-truth label for the unlabeled image, a pseudo-label for the weakly perturbed prediction is obtained by: $y^{w}\text{= argmax}(p^{w})$. Then, this pseudo-label is used to supervise the strongly perturbed prediction, as depicted in the standard flow presented in Fig. 1e (red arrow). This consistency term can be formulated as:
    $L_{u(standard)}=\sum\mathbb{1}(max(p^{w})\geq\tau)H(y^{w},p^{s}),$ (3)
    where $\mathbb{1}(\cdot)$ is an indicator function, and $\tau$ is a confidence threshold used to ignore noise in a pseudo-label.
  3. 3)
    Unsupervised – Proposed Flow To obtain additional pseudo-label guidance, the same weakly perturbed image, $x^{w}$, used in the standard branch, is fed into the previous model branch. Subsequently, $k$ models, where $k$ is chosen randomly from $\{1,2,...,K\}$, are selected from the saved list of previous models, and $k$ predictions for $x^{w}$ are generated. Previous guidance, $p^{w^{\prime}}$, is obtained by aggregating the $k$ predictions using randomized ensemble weights to improve the diversity of the pseudo-label view. As illustrated by the blue arrow in Fig. 1e, the previous guidance functions as additional regularizers, supervising the strongly perturbed prediction. This can be formulated as follows.
    $L_{u(prev)}=\sum\mathbb{1}(max(p^{w^{\prime}})\geq\tau)H(y^{w^{\prime}},p^{s}).$ (4)
@@
3.2 Previous Guidance